Three royal pythons and two anacondas measure 161 feet. The royal python's length increased by triple the anaconda's length is 119 feet. Find the maximum length for each snake.
Answer by mananth(16949)

Three royal pythons and two anacondas measure 161 feet. The royal python's length increased by triple the anaconda's length is 119 feet. Find the maximum length for each snake.
let royal python's length = x
Anaconda's length = y
3x+2y=161......................1
...
x+3y= 119.......................2
...
multiply (20 by -3
-3x-9y=-357
add to (1)
-7y=-357+161
-7y=-196
/-7
y= 28 feet Anaconda's lengtn
plug y in (1)
3x+2*28=161
3x+56=161
3x=161-56
3x=105
/3
x= 35 feet .... python's length